OpenStudy (anonymous):

A line passes through the point (0, –4) and has a slope of 3/2 What is the equation of the line?

OpenStudy (anonymous):

you can do point slope form which is y-k=m(x-h) where the point is (h,k) and m= slope so the equation would be y+4=3/2(x)

OpenStudy (anonymous):

thanks.